OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 1,024 and a4 = -16?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (tkhunny):

First Term = a = 1024 Second Term = ar = ?? Third Term = ar^2 = ?? Fourth Term = ar^3 = -16 Can you solve for r?

OpenStudy (anonymous):

umm im not sure how to

OpenStudy (tkhunny):

It's time to get sure. You have: a = 1024 ar^3 = -16 Solve for r. It may help to observe that 1024 = 2^10 and 16 = 2^4

OpenStudy (anonymous):

1073741840

OpenStudy (tkhunny):

Please demonstrate how you managed that. a = 1024 ar^3 = -16 Thus, but substitution 1024r^3 = -16 Thus, by division r^3 = -16/1024 = -(2^4)/(2^10) = -1/(2^6) Are we getting anywhere?

OpenStudy (anonymous):

1024^3=16 then i simplftyed we are getting somewhere

OpenStudy (tkhunny):

I see. You should have substitute ONLY for 'a'. The 'r' should not have gone anywhere. ar^3 = (a)*(r^3) Substituting for 'a' gives ar^3 = (a)*(r^3) = 1024(r^3)

OpenStudy (anonymous):

oh oj=k

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont know how to solve it

OpenStudy (tkhunny):

r^3 = -1/(2^6) You need to know how to solve that. Give it a go.

OpenStudy (anonymous):

r^3=-64

OpenStudy (tkhunny):

How did 2^6 magically get into the numerator? What you have shown, so far, is not encouraging. Have you been away from algebra for a while? You may need to consider backing up and doing a little more time in an earlier course. \(r^3 = -1/(2^6)\) \(\left(r^3\right)^{1/3} = \left(-1/(2^6)\right)^{1/3}\) \(r^{3\cdot(1/3)} = (-1)^{1/3}/(2^6)^{1/3}\) \(r = -1/\left(2^{6\cdot (1/3)}\right)\) \(r = -1/\left(2^{2}\right) = -1/4\) There should not be anything in there that you cannot do. If there is, please give that earlier course some thought.